OpenStudy (anonymous):

Hiiii guys :) I need help again with my math. Okay here's the question: Suppose you invest $2000 at an annual interest rate of 5.1% compounded continuously. How much will you have in the account after 3 years? Round the solution to the nearest dollar. I know how to start the equation but I always get the answer wrong no matter what. D: The formula is Pe^rt.

jimthompson5910 (jim_thompson5910):

Can you post what you've got so far?

OpenStudy (anonymous):

P=2000 r=0.051 t=3 A=Pe^rt A=2000*e^0.051*3 =2330 But apparently this is incorrect so. :(

jimthompson5910 (jim_thompson5910):

what did you get before you rounded?

OpenStudy (anonymous):

Oh right the rounding up. I got so frustrated with the problem I forgot about it. When I put it in the calculator it is 2330.649958. So since you round it up it would be 2331?

jimthompson5910 (jim_thompson5910):

correct, the answer is 2331

OpenStudy (anonymous):

WOOT! Thanks for your help! :D -hugs-

jimthompson5910 (jim_thompson5910):

you're welcome